# Data relativistic uncertainty framework for low-illumination anime scenery image enhancement

### v1

@@ 4.1 Implementation Details
Regarding the DRU framework, we implement the probability network $F_{q}$ by six choices of modern backbone networks including ResNet18, ResNet50 [he2016deep], MobileNetV2 [sandler2018mobilenetv2], VGG13BN, VGG19 [simonyan2014very] and ViT-B16 [dosovitskiy2020image]. These backbone networks follow the identical training data and recipe of the classifier $F_{dark/bright}$ as described in Sec. 3.1. The training recipe uses the pre-trained weights from ImageNet [deng2009imagenet], a batch size of 64 to run over 80 epochs, and the optimizer SGD [robbins1951stochastic] with learning rate of $1e^{-3}$ and momentum of 0.9. After obtaining the trained backbone networks to quantify the illumination uncertainty, we employ EnlightenGAN as the base architecture while retaining its default training hyperparameters [jiang2021enlightengan], and train six versions of EnlightenGAN using the DRU framework. We denote each version as DRU-EnlightenGAN-Backbone in the experiments. Our models are compared with five state-of-the-art (SOTA) unsupervised LLIE methods, such as Vanilla EnlightenGAN [jiang2021enlightengan], SCI [ma2022toward], ZeroDCE++ [li2021learning], RUAS [liu2021retinex] and CLIP-LIT [liang2023iterative]. To avoid the domain gap problem and ensure fair comparisons, trainDark and trainBright of the unpaired anime scenery dataset are adopted for all the experimental trainings. Specifically, our models, Vanilla EnlightenGAN and CLIP-LIT use trainDark and trainBright as unpaired dark-bright training sets, whereas, ZeroDCE++ combines trainDark and trainBright into a multi-exposure training set. SCI and RUAS use trainDark but not trainBright since they only need low-light training images. The training settings of SOTA methods follow the original practices of [jiang2021enlightengan, ma2022toward, li2021learning, liu2021retinex, liang2023iterative]. Lastly, testDark in the unpaired anime scenery dataset is utilized to evaluate the enhancement performance of trained models and the test results are concretely measured by four indicators involving BRISQUE [mittal2012no], PIQE [venkatanath2015blind], PI [blau20182018] and NIMA [talebi2018nima]. BRISQUE and PIQE assess the perceptual quality relative to the distortion level. PI measures the visual quality along with the reconstruction accuracy. NIMA estimates the aesthetic quality.
@@ 4.2 State-of-the-Art Comparison
Qualitative analysis. We compare the qualitative results of DRU-EnlightenGAN-ResNet18 with other SOTA methods as shown in Fig. 4. SCI, ZeroDCE++ and CLIP-LIT present inadequate results with varying degrees of dark coverage. ZeroDCE++ looks slightly better than the other two, but exists a gray color bias. Although the results of RUAS are brighter than all other methods, they suffer from overexposed artifacts, leading to a noticeable blue color cast. Last two rows exhibit more pleasing visual results than their counterparts and furthermost, DRU-EnlightenGAN-ResNet18 outperforms Vanilla EnlightenGAN with less shadow and color bias. Here, Fig. 4(row,column) is used to denote the image position. The local region between boat and grass in Fig. 4(6,1) displays darker and larger shadow than that of Fig. 4(7,1). The similar phenomena occur on the pillar in Fig. 4(6,2) and Fig. 4(7,2), the left part of lake in Fig. 4(6,3) and Fig. 4(7,3), and the right region of sky in Fig. 4(6,5) and Fig. 4(7,5). Besides that, the tree of Fig. 4(6,4) shows a yellow cast while that of Fig. 4(7,4) is naturally close to green. Both of them preserve clearer foreground textures and more complete background details compared with other methods. Therefore, the above visual analysis sufficiently proves the effectiveness of DRU framework.
@@
Quantitative analysis. By referring to Table 1, the DRU versions of EnlightenGAN basically surpass all the SOTA counterparts across four indicators by a large margin, which further demonstrates the superiority and practicality of DRU framework. The aim of low-illumination enhancement for anime scenery images is not simply to improve or raise the brightness. Even though RUAS attains the brightest results in Fig. 4, it performs the bottom in all metrics, which reveals that RUAS is unable to learn a rational illumination concept from the low-light images in trainDark, aligning with the discovery in [wang2024zero]. Note that Vanilla EnlightenGAN achieves better scores than other SOTA methods, followed by SCI. For simplicity, we refer to each DRU version of EnlightenGAN by its backbone below. ResNet50 and ResNet18 shows the first and second score for BRISQUE. ResNet18 and MobileNetV2 are ranked as the first and second for PIQE. ViT-B16 is the only one that reaches the best scores for PI and NIMA, respectively followed by ResNet50 and VGG13BN. Consequently, DRU-EnlightenGAN-ViT-B16 stands out among all the methods in terms of overall quality scores. This implies that the Transformer structure can exert a stronger generalization capacity than CNN in quantifying the illumination uncertainty.
@@ 4.3 Ablation Study
In this section, we use DRU-EnlightenGAN-ResNet18 and Vanilla EnlightenGAN for the ablation experiments w.r.t two crucial questions, where they are shortly denoted as DRU and Vanilla.
Are both confident and uncertain dark/bright images necessary for DRU training? To examine the effect of confident and uncertain images to DRU and Vanilla training respectively, we separate the confident and uncertain images from trainDark and trainBright, which are denoted as Confident and Uncertain for training. So Confident+Uncertain is equivalent to all the images of trainDark and trainBright. The goal of enhancement task takes into account both perceptual and aesthetic qualities, namely, the results for BRISQUE and NIMA are expected to approach lower and higher respectively. From this view, DRU on Confident+Uncertain demonstrates the best trade-off between the perceptual and aesthetic qualities in Table 2. With only Confident, DRU performs well in BRISQUE but poorly in NIMA, indicating that the enhanced results are bright but not smooth and aesthetically pleasing. However, using uncertainty only for DRU leads to relatively dark but smooth enhanced results. Thus, both confident and uncertain images are needed for DRU training since confident images are strong priors for DRU to learn towards better perceptual quality while uncertain images benefit DRU from the illumination uncertainty to learn greater aesthetic quality. Although the overall performance of Vanilla is suboptimal to DRU. Interestingly, we observe from Table 2 that Vanilla obtains better BRISQUE and NIMA scores on Uncertain than another two settings, followed by Confident+Uncertain. This indicates the importance of uncertain images for Vanilla training and the necessity of DRU to exploit illumination uncertainty of confident and uncertain images for further learning.
@@
We reclassify the original trainDark and trainBright using an unreliable pretrained ResNet18 to produce new trainDark and trainBright with severe data noise. Table 3 shows the performance of DRU and Vanilla after training on Original and Noisy sets. Between the Original and Noisy sets, the BRISQUE difference of DRU is 0.8565, which is much smaller than the 1.8664 of Vanilla. The NIMA difference of DRU is 0.0394, less than the 0.056 of Vanilla. Vanilla trained on Noisy produces darker enhanced results that are given the higher BRISQUE score, and darker images tend to exhibit more smooth textures, affecting its NIMA score higher. So this signifies worse performance. DRU achieves better trade-off between BRISQUE and NIMA scores and less variation for each metric. Hence, DRU is able to reduce the performance loss raised by data noise due to misclassification compared to Vanilla.
@@ 5 Conclusion
We address the domain gap and data shortage problems faced by low-illumination enhancement for anime scenery images through establishing the first unpaired anime scenery dataset. In addition, we propose a data relativistic uncertainty framework which enables the enhancement task to learn under data uncertainty perspective. Our framework is applicable not only to EnlightenGAN, but also to any state-of-the-art methods as long as their loss terms can be adapted to use the framework. Experimental results evince its advantageous performance across multiple quality metrics. Our framework provides the possibility for the realization of future work in [jiang2021enlightengan]. We also expect that this new paradigm of data uncertainty learning can generalize to other vision and language tasks.

### v2

@@ 4.1 Implementation Details
Regarding the DRU framework, we implement the probability network $F_{q}$ by six choices of modern backbone networks including ResNet18, ResNet50 [14], MobileNetV2 [42], VGG13BN, VGG19 [45] and ViT-B16 [9]. These backbone networks follow the identical training data and recipe of the classifier $F_{dark/bright}$ as described in Sec. 3.1. The training recipe uses the pre-trained weights from ImageNet [8], a batch size of 64 to run over 80 epochs, and the optimizer SGD [41] with learning rate of $1e^{-3}$ and momentum of 0.9.
After obtaining the trained backbone networks to quantify the illumination uncertainty, we employ EnlightenGAN as the base architecture while retaining its default training hyperparameters [17], and train six versions of EnlightenGAN using the DRU framework. We denote each version as DRU-EnlightenGAN-Backbone in the experiments. Our models are compared with five state-of-the-art (SOTA) unsupervised LLIE methods, such as Vanilla EnlightenGAN [17], SCI [32], ZeroDCE++ [25], RUAS [28] and CLIP-LIT [27].
To avoid the domain gap problem and ensure fair comparisons, trainDark and trainBright of the unpaired anime scenery dataset are adopted for all the experimental trainings. Specifically, our models, Vanilla EnlightenGAN and CLIP-LIT use trainDark and trainBright as unpaired dark-bright training sets, whereas, ZeroDCE++ combines trainDark and trainBright into a multi-exposure training set. SCI and RUAS use trainDark but not trainBright since they only need low-light training images. The training settings of SOTA methods follow the original practices of [17, 32, 25, 28, 27]. Lastly, testDark in the unpaired anime scenery dataset is utilized to evaluate the enhancement performance of trained models and the test results are concretely measured by four indicators involving BRISQUE [36], PIQE [49], PI [1] and NIMA [47]. BRISQUE and PIQE assess the perceptual quality relative to the distortion level. PI measures the visual quality along with the reconstruction accuracy. NIMA estimates the aesthetic quality. Additionally, we conduct a user study to evaluate the perceptual and aesthetic quality of different models.
@@ 4.2 State-of-the-Art Comparison
Qualitative analysis. We compare the qualitative results of DRU-EnlightenGAN-ResNet18 with other SOTA methods as shown in Fig. 4. SCI, ZeroDCE++ and CLIP-LIT present inadequate results with varying degrees of dark coverage. ZeroDCE++ looks slightly better than the other two, but exists a gray color bias. Although the results of RUAS are brighter than all other methods, they suffer from overexposed artifacts, leading to a noticeable blue color cast. Last two rows exhibit more pleasing visual results than their counterparts and furthermost, DRU-EnlightenGAN-ResNet18 outperforms Vanilla EnlightenGAN with less shadow and color bias. Here, Fig. 4(row,column) is used to denote the image position.
The local region between boat and grass in Fig. 4(6,1) displays darker and larger shadow than that of Fig. 4(7,1). The similar phenomena occur on the pillar in Fig. 4(6,2) and Fig. 4(7,2), the left part of lake in Fig. 4(6,3) and Fig. 4(7,3), and the right region of sky in Fig. 4(6,5) and Fig. 4(7,5). Besides that, the tree of Fig. 4(6,4) shows a yellow cast while that of Fig. 4(7,4) is naturally close to green. Both of them preserve clearer foreground textures and more complete background details compared with other methods. Therefore, the above visual analysis sufficiently proves the effectiveness of DRU framework.
@@
Quantitative analysis. By referring to Table 1, the DRU versions of EnlightenGAN basically surpass all the SOTA counterparts across four indicators by a large margin, which further demonstrates the superiority and practicality of DRU framework. The aim of low-illumination enhancement for anime scenery images is not simply to improve or raise the brightness. Even though RUAS attains the brightest results in Fig. 4, it performs the bottom in all metrics, which reveals that RUAS is unable to learn a rational illumination concept from the low-light images in trainDark, aligning with the discovery in [51]. Note that Vanilla EnlightenGAN achieves better scores than other SOTA methods, followed by SCI.
For simplicity, we refer to each DRU version of EnlightenGAN by its backbone below. ResNet50 and ResNet18 shows the first and second score for BRISQUE. ResNet18 and MobileNetV2 are ranked as the first and second for PIQE. ViT-B16 is the only one that reaches the best scores for PI and NIMA, respectively followed by ResNet50 and VGG13BN.
Consequently, DRU-EnlightenGAN-ViT-B16 stands out among all the methods in terms of overall quality scores. This implies that the Transformer structure can exert a stronger generalization capacity than CNN in quantifying the illumination uncertainty.
Table 2: The user study results are reported as average preference scores (APS), representing the proportion of times each method is selected as the best. Bold indicates the best score, and underline indicates the second best.
Method APS (%)
SCI 6.5650
RUAS 14.3770
ZeroDCE++ 6.5650
CLIP-LIT 5.0030
Vanilla EnlightenGAN 33.1270
DRU-EnlightenGAN-ResNet18 34.3790
User study. We conduct a user study with 32 participants to evaluate the visual quality of anime scenes produced by each method. Participants select their preferred options from six candidates across 10 image sets, according to three criteria: appropriate brightness, naturalness of colors, and adherence to anime aesthetics. As summarized in Table 2, the average preference scores show that DRU-EnlightenGAN-ResNet18 outperforms all other methods, demonstrating the effectiveness of our framework in refining EnlightenGAN.
@@ 4.3 Ablation Study
In this section, two initial ablation studies are conducted using DRU-EnlightenGAN-ResNet18 and Vanilla EnlightenGAN, hereafter referred to as DRU and Vanilla, to investigate two key questions. Details of the remaining experiments are provided in the following sections.
Are both confident and uncertain dark/bright images necessary for DRU training? To examine the effect of confident and uncertain images to DRU and Vanilla training respectively, we separate the confident and uncertain images from trainDark and trainBright, which are denoted as Confident and Uncertain for training. So Confident+Uncertain is equivalent to all the images of trainDark and trainBright. The goal of enhancement task takes into account both perceptual and aesthetic qualities, namely, the results for BRISQUE and NIMA are expected to approach lower and higher respectively. From this view, DRU on Confident+Uncertain demonstrates the best trade-off between the perceptual and aesthetic qualities in Table 3.
With only Confident, DRU performs well in BRISQUE but poorly in NIMA, indicating that the enhanced results are bright but not smooth and aesthetically pleasing. However, using uncertainty only for DRU leads to relatively dark but smooth enhanced results. Thus, both confident and uncertain images are needed for DRU training since confident images are strong priors for DRU to learn towards better perceptual quality while uncertain images benefit DRU from the illumination uncertainty to learn greater aesthetic quality.
Although the overall performance of Vanilla is suboptimal to DRU. Interestingly, we observe from Table 3 that Vanilla obtains better BRISQUE and NIMA scores on Uncertain than another two settings, followed by Confident+Uncertain. This indicates the importance of uncertain images for Vanilla training and the necessity of DRU to exploit illumination uncertainty of confident and uncertain images for further learning.
@@
We reclassify the original trainDark and trainBright using an unreliable pretrained ResNet18 to produce new trainDark and trainBright with severe data noise. Table 4 shows the performance of DRU and Vanilla after training on Original and Noisy sets.
Between the Original and Noisy sets, the BRISQUE difference of DRU is 0.8565, which is much smaller than the 1.8664 of Vanilla. The NIMA difference of DRU is 0.0394, less than the 0.056 of Vanilla. Vanilla trained on Noisy produces darker enhanced results that are given the higher BRISQUE score, and darker images tend to exhibit more smooth textures, affecting its NIMA score higher. So this signifies worse performance. DRU achieves better trade-off between BRISQUE and NIMA scores and less variation for each metric. Hence, DRU is able to reduce the performance loss raised by data noise due to misclassification compared to Vanilla.
Table 5: Performance of different methods on the unseen dataset. Input denotes the test images. We report the scores as BRISQUE↓||NIMA↑. Vanilla and DRU-ViT-B16 correspond to two different versions of EnlightenGAN. Bold marks the best score, underline marks the second best.
Method BRISQUE↓||NIMA↑
Input 37.8164 $|$ 3.9054
SCI 31.6215 $|$ 3.8296
RUAS 35.4778 $|$ 3.5323
ZeroDCE++ 33.9296 $|$ 3.6256
CLIP-LIT 33.1846 $|$ 3.8368
Vanilla 32.0589 $|$ 3.9199
DRU-ViT-B16 31.1713 $|$ 3.9299
Evaluation on unseen dataset. To further evaluate the generalizability of different methods, we construct an unseen dark dataset of 738 anime scenery images sourced from [59, 7, 22]. As shown in Table 5, DRU-ViT-B16 obtains the best performance among all methods, indicating robust perceptual and aesthetic generalization on unseen data.
Table 6: Performance of RUAS variants on testDark and unseen datasets. Scores are reported as BRISQUE↓||NIMA↑. RUAS is the vanilla variant. RUAS-DRU denotes the RUAS variant trained with DRU-ResNet18. Bold indicates the best score.
Variant testDark unseen
RUAS 49.1256 $|$ 4.1766 35.4778 $|$ 3.5323
RUAS-DRU 34.6225 $|$ 4.2826 35.4175 $|$ 3.5390
Transferability of DRU. To assess the generality of our framework, we adapt DRU to train a variant of RUAS. Table 6 compares the performance of RUAS and RUAS-DRU on testDark and the unseen dataset. Training RUAS with DRU yields notable improvements in both perceptual and aesthetic quality, highlighting the framework’s effectiveness and its potential applicability to low-light enhancement beyond EnlightenGAN.
@@ 5 Conclusion
We address the domain gap and data shortage problems faced by low-illumination enhancement for anime scenery images through establishing the first unpaired anime scenery dataset. In addition, we propose a Data Relativistic Uncertainty (DRU) framework that reformulates low-light enhancement learning from a data uncertainty perspective, enabling training with uncertain samples rather than relying solely on confident ones. Our framework is applicable not only to EnlightenGAN, but also to any state-of-the-art methods as long as their loss terms can be adapted to use the framework. Experimental results evince its advantageous performance across multiple quality metrics. Our framework provides the possibility for the realization of future work in [17]. We also expect that this new paradigm of data uncertainty learning can generalize to other vision and language tasks.